Suppose that Bob's knapsack private key is (3, 5, 10, 23), multiplier is m-1 = 6, and modulus n = 47. What is the plaintext (in hex) if the ciphertext C = 20?
Answer #1

Solution:

given m-1=6 modulus n=47 ciphertext=20 private key ={3,5,10,23}

plaintext= ciphertext   m-1 mod n. Therefore, 20 6 mod 47= 26

Now we have to find out by adding which keys from the set of private keys can we form 26, the just place 1 for those digits and 0 for others.

we know 23+3=26 therefore 1001. so the plain text is 1001.

Please give thumbsup, if you like it. Thanks.